3x-y=18
x+2y=13

Solve use elimination method

Answer by oscargut(2103)

3x-y=18
x+2y=13

If we multiply the first equation by 2:
6x-2y=36
x+2y=13
Adding:
7x = 49
x = 7
y = 3
You can ask me more at: mthman@gmail.com
Thanks